OpenStudy (anonymous):

Describe the differences between the graph of y equals negative 3 over 4 open parentheses x minus 6 close parentheses squared plus 4 and the standard position graph of y = x2. please help

OpenStudy (ranga):

\[y = -\frac{ 3 }{ 4}(x-6)^2 + 4 \quad \text{compared with} \quad y = x^2\]

OpenStudy (ranga):

Replacing x in the equation y = x^2 with (x-6) is equivalent to shifting the graph y = x^2 to the RIGHT by 6 units. Changing y = x^2 to y = x^2 + 4 is equivalent to shifting the graph UP by 4 units. continued...

OpenStudy (ranga):

In y = x^2, the coefficient of x^2 is positive. Therefore, the graph, which is a parabola, opens upwards. If the coefficient is negative, the graph will be an inverted parabola which will open downward. So y = -x^2 is a parabola that opens downward. If the -1 coefficient were changed to a fraction, such as -3/4, the graph will be compressed vertically (a little bit squished).

OpenStudy (ranga):

So starting with y = x^2 you can get the other graph by shifting it to the right by 6 units, lifting it up by 4 units, flipping it so it opens downward and squishing it slightly. y = x^2 has vertex at (0,0). It has a minimum at the vertex and the minimum value is 0. It is a parabola that opens upward. It touches the x-axis at just one point. y = -3/4(x-6)^2 + 4 has the vertex at (6,4). Because it is an inverted parabola, it attains its maximum at x = 6 and the maximum value is 4. It cuts the x axis at two points.